OpenStudy (anonymous):

Kate's father is three times her age. How old are Kate and her father now if, in 4 years time, the sum of their ages is 68? Solve and show all working out.

OpenStudy (thesecret20111):

Try to write an equation for this it'll make it very easy

OpenStudy (anonymous):

15*3=45 45+15=60 60+8=68